How many 3 digit counting numbers are even?
Found 2 solutions by fractalier, AnlytcPhil:
Answer by fractalier(6550)

Well, half of ALL counting numbers are even, so if we can figure out how many three-digit counting numbers there are, we are set...
They start at 100 and end at 999, a total of 900 numbers.
Thus 450 of them are even.

Answer by AnlytcPhil(1810)

How many 3 digit counting numbers are even?

This is the arithmetic sequence

100, 102, 104, ··· , 998

an = a1 + (n-1)d

To find d:

d = 102 - 100 = 104 - 102 = 2

a1 = 100

an = 998

an = a1 + (n-1)d

998 = 100 + (n-1)(2)

998 = 100 + 2(n-1)

998 = 100 + 2n - 2

998 = 98 + 2n

900 = 2n

450 = n

So there are 450 three-digit counting numbers

Edwin
AnlytcPhil@aol.com